OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form. the square root of 169 − the square root of 50 − the square root of 8?

OpenStudy (anonymous):

13^2 = 169, so Sqrt[169] = 13 50 = 2 * 25, 5^2=25, so Sqrt[50] = 5Sqrt[2] 8 = 2*4, 2^2=4, so Sqrt[8] = 2Sqrt[2]

OpenStudy (anonymous):

thank you william. briand:D